Quantrex Quantrex AcademyJEE · NEET · NDA PYQs with solutions Open app
NDA 2023
Exam
NDA
Year
2023
Subject
Mathematics
Chapter
Probability

NDA 2023 · Mathematics · Probability

Let A and B be two independent events such that P( A )=0.7, P( B...

Let AA and BB be two independent events such that P(Aˉ)=0.7,P(Bˉ)=k,P(A∪B)=0.8P(\bar{A})=0.7, P(\bar{B})=k, P(A \cup B)=0.8 What is the value of k ?

Options

  1. A57\frac{5}{7}
  2. B00
  3. C27\frac{2}{7}
  4. D17\frac{1}{7}

Correct answer

C. 27\frac{2}{7}

Step-by-step solution

P(Aˉ)=0.7⇒P(A)=0.3P(Bˉ)=k⇒P(B)=1−kP(A∪B)=0.8P(\bar{A})=0.7 \Rightarrow P(A)=0.3P(\bar{B})=k \Rightarrow P(B)=1-kP(A \cup B)=0.8 For independent events P(A∩B)=P(A)⋅P(B)P(A∪B)=P(A)+P(B)−P(A)⋅P(B)⇒0.8=0.3+1−k−0.3(1−k)⇒0.5=1−k−0.3+0.3k⇒0.5=0.7−0.7k⇒0.7k=0.2⇒k=27P(A \cap B)=P(A) \cdot P(B)P(A \cup B)=P(A)+P(B)-P(A) \cdot P(B)\Rightarrow \quad 0.8=0.3+1-k-0.3(1-k)\Rightarrow \quad 0.5=1-k-0.3+0.3 k\Rightarrow \quad 0.5=0.7-0.7 k\Rightarrow \quad 0.7 k=0.2 \Rightarrow k=\frac{2}{7}
Practice Probability on Quantrex Academy →

More from Probability

Passage: Four dice are rolled. Question: What is the probability that getting a total of… NDA 2026Three events A , B and C are such that A and B are disjoint, A and C are independent, B… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026Passage: Four dice are rolled. Question: What is the probability that getting a total of… NDA 2026For a Binomial distribution with mean 6 and standard deviation 2 , what is the value of… NDA 2026Passage: A 4-digit number is selected at random formed by using the digits 0 , 1 , 2 , 3… NDA 2026In an entrance test there are multiple choice questions. There are four options for each… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026Consider the following statements in respect of the events A , B , C : I. (A B C) ( A B C… NDA 2026Passage: Let A , B , C and D be mutually exclusive and exhaustive events and… NDA 2026The probability that a family owns a laptop is 0.68 ; that it also owns a desktop is 0.56… NDA 2025An urn contains 10 white and 5 red balls. If two balls are drawn at random, then what is… NDA 2025An urn contains 5 white, 6 red and 4 blue balls. Three balls are drawn at random. What is… NDA 2025Under which of the following conditions may binomial distribution be used? I. The number… NDA 2025Three faces of a die are black, two faces are white and one face is red. The die is… NDA 2025A person X speaks the truth 4 out of 5 times and person Y speaks the truth 5 out of 6… NDA 2025 Full Probability list Probability JEE Maths PYQsJEE Mathematics PYQs All NDA PYQs